Campuses:
Link to main foregrounds telecons page
Link to logbook
Attendance: Ragnhild, Charles, Mathieu, Sebastian, Julien, Jacques
Regrets:
Commander - Good progress. At the previous telecon the r-estimations seemed biased high. Ragnhild optimized priors. B-mode spectra show no clear indication of significant excess power at low ell on B-mode Cl plots. New priors reduce a bit the power at high galactic latitude. (note: the changes were mean value of dust temperature 21—> 23K, beta: 1.5 —> 1.55).
NILC - Jacques still has to produce masks and send them to Mathieu for model 98. - No new results (Mathieu still tying to run his codes on the spanish cluster). - Mathieu to provide the canonical mask he is using to the delensing team. A slightly larger sky fraction will be used for delensing to avoid edge effects on the useful sky region.
DELENSING - Still waiting for the nside=2048 outputs of NILC. - For the moment Mathieu will put at NERSC the 512 outputs so that the pipeline can be set-up by Sebastian and Julien. - Need residual noise maps, Mathieu will make the 10 realizations available at NERSC.
TEXT - postponed… it is not the most urgent part of the work.